Hi could you help me to understand how I can simplify: 2(4x^2+7x)-(x+2)(8x-2)
Thank you
1
Expert's answer
2012-02-03T07:11:41-0500
& First let's open the braces:
2(4x^2+7x)-(x+2)(8x-2)=8x^2+14x-(8x^2-2x+16x-4)=8x^2+14x-8x^2-14x+4=4 so& no further actions are required, we got single number
Remember, when you open the braces before which there is "-" sign you should change the sigh of every term inside the braces when opening them.